# Semantics-Oriented Multitask Learning for DeepFake Detection: A Joint Embedding Approach

### v1

@@ Abstract
In recent years, the multimedia forensics and security community has seen remarkable progress in multitask learning for DeepFake (i.e., face forgery) detection. The prevailing strategy has been to frame DeepFake detection as a binary classification problem augmented by manipulation-oriented auxiliary tasks. This strategy focuses on learning features specific to face manipulations, which exhibit limited generalizability. In this paper, we delve deeper into semantics-oriented multitask learning for DeepFake detection, leveraging the relationships among face semantics via joint embedding. We first propose an automatic dataset expansion technique that broadens current face forgery datasets to support semantics-oriented DeepFake detection tasks at both the global face attribute and local face region levels. Furthermore, we resort to joint embedding of face images and their corresponding labels (depicted by textual descriptions) for prediction. This approach eliminates the need for manually setting task-agnostic and task-specific parameters typically required when predicting labels directly from images. In addition, we employ a bi-level optimization strategy to dynamically balance the fidelity loss weightings of various tasks, making the training process fully automated. Extensive experiments on six DeepFake datasets show that our method improves the generalizability of DeepFake detection and, meanwhile, renders some degree of model interpretation by providing human-understandable explanations.
@@ I Introduction
The emergence of deep learning models [1, 2, 3, 4] has greatly simplified and automated the production of realistic counterfeit face images and videos, commonly referred to as DeepFakes. This seriously threatens the authenticity and integrity of digital visual media. In response, a wide array of DeepFake detection methods have been developed to efficiently screen manipulated face imagery.
@@
The early DeepFake detectors [5, 6, 7] were largely shaped by established photo forensics techniques [8], through the analysis of statistical anomalies [9], visual artifacts [10, 11, 12], and physical inconsistencies [13, 14, 15]. The advent of deep learning has spurred the development of numerous learning-based detection methods [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37], which operate as binary classifiers to predict whether a test image is real or fake. This formulation can be augmented by incorporating auxiliary tasks, such as manipulation type identification [38, 39], manipulation parameter estimation [40], blending boundary detection [21], and face reconstruction [23, 26]), leading to a multitask learning paradigm [20, 21, 22, 23, 24, 25, 26, 27]. The underlying assumption is that the primary task will likely gain from including auxiliary tasks through knowledge transfer. However, most auxiliary tasks are specific to manipulations, which may hinder the ability to generalize. Recently, Zou et al. [41] introduced a semantics-oriented multitask learning paradigm for DeepFake detection. They defined face forgery as the use of computational methods that modify semantic face attributes beyond the thresholds of human discrimination. Although conceptually and computationally appealing, training semantics-oriented DeepFake detectors would require considerable human annotations to specify manipulation degree parameters and semantic label hierarchies [41]. Moreover, the allocation of task-agnostic and task-specific parameters is optimized manually in multitask learning, which is bound to be suboptimal in facilitating knowledge transfer among tasks.
Figure 1: Illustration of the label hierarchy of the expanded FF++ dataset [18] by the proposed semantics-oriented dataset expansion technique.
@@
In this paper, we further pursue the promising direction of semantics-oriented multitask learning for DeepFake detection with two important improvements. First, we present a dataset expansion technique that automatically enriches the binary labeling of existing DeepFake datasets to semantic hierarchical labeling while increasing their sizes (see Fig. 1). A crucial element of our technique is to incorporate the physical_inconsistency node that encompasses local face region manipulations, without manually identifying interdependent changes in global face attributes (if any). Second, unlike discriminative learning that directly predicts target labels from images, we take inspiration from the recent advances in vision-language correspondence [42] and encode both the input image and the target label hierarchy (represented by text templates) into a shared feature space. This enables us to view model parameter sharing/splitting in multitask learning from the new perspective of model capacity allocation. That is, we share the image encoder parameters across all tasks, with the capacity allocated to individual tasks being end-to-end optimized. In addition, we adopt bi-level optimization [43] to minimize the primary task with respect to the fidelity loss weightings [44] in the validation set as a way of task prioritization. The parameters of the DeepFake detector are learned by minimizing the weighted summation of fidelity losses for individual tasks.
@@
We demonstrate the effectiveness of our Semantics-oriented Joint Embedding DeepFake Detector (SJEDD) through extensive experiments on six datasets [18, 45, 46, 47, 48, 41]. SJEDD consistently improves the cross-dataset, cross-manipulation, and cross-attribute detection performance compared to eighteen state-of-the-art DeepFake detectors.
@@ II Related Work
In this section, we provide a concise review of DeepFake detectors, emphasizing the multitask learning approach. We also discuss joint embedding architectures in the broader context of machine learning.
@@ II-A DeepFake Detectors
Non-learning-based detectors rely on extensive human expertise for their design. The first type of knowledge involves mathematical models of the image source, which summarizes the essential elements of what constitutes an authentic image [9]. The second type of knowledge involves computational models of the digital photography pipeline, along which each step can be exploited to expose forgeries [49], including the lens distortion and flare [10, 11], color filter array [50], sensor noise [12], pixel response non-uniformity [51], camera response function [52], and file storage format [53]. The third type of knowledge involves physical and geometric models of the three-dimensional world, especially those that capture the interactions of light, optics, and objects. The image forgeries can be exposed by detecting physical and geometric inconsistencies, such as those found in illumination [13], shadow [14], reflection [15], and motion [54]. Generally, non-learning-based methods are limited by the precision and expressiveness of manually crafted models and thus achieve limited success.
@@
Learning-based methods detect DeepFake images and videos by learning directly from data. Afchar et al. [17] proposed to examine DeepFake videos at the mesoscopic level. Zhou et al. [16] designed a two-stream network to jointly analyze camera characteristics and assess visual artifacts. Li and Lyu [55] aimed to detect affine warping artifacts, while Li et al. [21] focused on locating image blending boundaries in the intermediate steps of face manipulations. Dong et al. [25] revealed a shortcut termed “implicit identity leakage,” and described an identity-unaware DeepFake detector. Gu et al. [56] and Zheng et al. [57] examined spatiotemporal inconsistencies at the signal level, while Haliassos et al. [28] did so at the semantic level through lipreading pretraining. Xu et al. [58] proposed a thumbnail layout to trade spatial modeling for temporal modeling. Another line of research that has received significant attention involves detecting abnormal physiological reactions from blinking of the eyes [59], pupil shape [60], head pose [61], corneal specularity [62], and action units [63]. Complex architectures, such as capsule networks [64] and vision Transformers [65, 31], along with advanced training strategies such as adversarial training [24], test-time training [66], graph learning [34, 36], contrastive learning [32, 37], and out-of-distribution detection [67], have also been used for DeepFake detection.
@@
Multitask learning-based methods have emerged as a promising offshoot of learning-based methods for DeepFake detection in the past five years. Notable auxiliary tasks include manipulation localization [20, 24] (with blending boundary detection [21] as a special case), manipulation classification and attribution [38, 39], manipulation parameter estimation [40], face image reconstruction [23, 26], and tasks that promote robust and discriminative feature learning [27]. Most auxiliary tasks are manipulation-oriented, which may increase the risk of overfitting and result in poor generalization to novel manipulations. This work improves upon the recently proposed semantics-oriented multitask learning paradigm [41]. From the data perspective, we introduce a dataset expansion technique that not only increases the sizes of existing DeepFake datasets but also enriches them with a collection of labels organized in a semantic hierarchical graph (see Fig. 1). From the model perspective, we leverage joint embedding instead of discriminative learning to automate model parameter sharing/splitting and to facilitate task relationship modeling. Concurrently, Sun et al. [33] proposed VLFFD, which also implemented DeepFake detection via joint embedding. The proposed SJEDD differs significantly from VLFFD in terms of motivation (i.e., enhancing semantics-oriented detection as opposed to leveraging multimodal data), task nature (i.e., semantics-oriented tasks rather than manipulation-oriented ones), and model optimization (i.e., principled bi-level optimization instead of heuristic coarse-to-fine training). Consequently, under the same experimental setups, the proposed SJEDD outperforms VLFFD on various DeepFake datasets by wide margins (see Sec. V-B).
@@
Figure 2: Pipeline for local face region manipulation. When searching for the best candidate as the source image, we minimize the Euclidean distance between $68$ detected landmarks [21] of the target and candidate face images (excluding those with the same identity).
@@ II-B Joint Embedding Architectures
Joint embedding architectures [68] aim to learn embeddings that are close for compatible inputs (e.g., an input image and its target label) and distant for incompatible inputs, as opposed to discriminative approaches that directly predict labels from inputs. The idea can be traced back to the Siamese network [69], initially proposed for signature verification and later adapted to computer vision applications like face verification [70] and recognition [71], and video feature learning [72]. With the rise of self-supervised learning, a thorough exploration of joint embedding architectures is conducted, contrastively (e.g., MoCo [73] and SimCLR [74]) and non-contrastively (e.g., Barlow Twins [75] and I-JEPA [76]). The more recent vision-language models [77, 42] are also manifestations of joint embedding architectures and are quickly adapted to multiple downstream tasks such as image quality assessment [78], realistic image enhancement [79], image-text retrieval [80], and multimodal fake news detection [81]. Here, we use CLIP [42], a vision-language model pre-trained on massive image-text pairs, to implement SJEDD.
@@ III Semantics-Oriented Dataset Expansion: A Demonstration on FF++
In this section, we describe our dataset expansion method in detail, consisting of two steps: data augmentation and data relabeling. Data augmentation generates new images by manipulating local face regions. Data relabeling automatically assigns DeepFake images with a set of labels through semantic contextualization of face manipulations [41]. To demonstrate the proposed dataset expansion technique, we apply it to expand the widely used FF++ dataset [18], while acknowledging that our method can also be applied to other DeepFake datasets.
@@ III-A Data Augmentation
We augment FF++ by manipulating four local face regions, i.e., eye, lip, mouth, and nose, as illustrated in Fig. 2. Given a real face as the target, we initially search for the best source face (real or fake) by minimizing the Euclidean distance between their corresponding $68$ detected landmarks [21] while excluding faces with the same identity [21]. Next, we linearly blend the source local region, indicated by a Gaussian-blurred region-of-interest mask, into the target image, followed by color correction [82].
III-B Data Relabeling
Inspired by [41], we first construct a semantic hierarchical graph to contextualize the four face manipulations in FF++: Deepfakes [1], FaceSwap [83], Face2Face [84], and NeuralTextures [4]. The former two manipulations swap faces between images to achieve identity manipulation, while the latter two manipulations change the expression of the target face through manipulation of the mouth region. Thus, we instantiate two global face attribute nodes: expression and identity. We also include six leaf nodes, representing nonoverlapping local face regions: eye, eyebrow, lip, mouth, nose, and skin. For the two identity manipulations based on face swapping, all local face regions are assumed to be altered. In contrast, for the two expression manipulations, only lip and mouth regions are affected. To accommodate local face manipulations in the graph, we introduce an additional global face attribute node—physical_inconsistency—which signifies images where a local face region is replaced by the corresponding region from another image (real or fake), resulting in physical incompatibility between this local region and the other real face regions.
TABLE I: Summary of the expanded FF++ dataset by the proposed dataset expansion technique, where “1” means the corresponding face attribute (or region) is modified and “0” otherwise. “N/A” indicates the manipulation status is not accessible (i.e., unobserved)
Manipulation / Region(s) Face Expression Identity Physical_inconsistency Eye Eyebrow Lip Mouth Nose Skin
@@ III-C Discussion
The proposed dataset expansion technique allows some labels in the semantic hierarchical graph to remain unobserved, which helps encourage labeling reliability. For instance, in the case of face swapping, it is unclear whether the expression has been altered, so we leave this label unobserved. Meanwhile, in the expanded dataset, a single face manipulation can simultaneously alter multiple face attributes (e.g., identity and physical_consistency by FaceSwap [83]), while different manipulations can also modify the same attribute (e.g., identity by Deepfakes [1] and FaceSwap). That is, our technique emphasizes a semantics-oriented rather than manipulation-oriented contextualization of face manipulations, akin to FFSC [41], but achieved without expensive human labeling. Table I presents the list of face manipulations in the expanded FF++ dataset, along with the corresponding modified face attributes and regions.

### v2

@@ Abstract
In recent years, the multimedia forensics and security community has seen remarkable progress in multitask learning for DeepFake (i.e., face forgery) detection. The prevailing approach has been to frame DeepFake detection as a binary classification problem augmented by manipulation-oriented auxiliary tasks. This scheme focuses on learning features specific to face manipulations with limited generalizability. In this paper, we delve deeper into semantics-oriented multitask learning for DeepFake detection, capturing the relationships among face semantics via joint embedding. We first propose an automated dataset expansion technique that broadens current face forgery datasets to support semantics-oriented DeepFake detection tasks at both the global face attribute and local face region levels. Furthermore, we resort to the joint embedding of face images and labels (depicted by text descriptions) for prediction. This approach eliminates the need for manually setting task-agnostic and task-specific parameters, which is typically required when predicting multiple labels directly from images. In addition, we employ bi-level optimization to dynamically balance the fidelity loss weightings of various tasks, making the training process fully automated. Extensive experiments on six DeepFake datasets show that our method improves the generalizability of DeepFake detection and renders some degree of model interpretation by providing human-understandable explanations.
@@ I Introduction
The emergence of deep learning [1, 2, 3, 4] has greatly simplified and automated the production of realistic counterfeit face images and videos, commonly referred to as DeepFakes. This seriously threatens the authenticity and integrity of digital visual media. In response, a wide array of DeepFake detection methods have been developed to efficiently screen manipulated face imagery [5, 6].
@@
Early DeepFake detectors [7, 8, 9] were largely shaped by established photo forensics techniques [10], through the analysis of statistical anomalies [11], visual artifacts [12, 13, 14], and physical inconsistencies [15, 16, 17]. The advent of deep learning has spurred the development of numerous learning-based detection methods [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38], operating as binary classifiers to predict whether a test image is real or fake. This formulation can be augmented by various auxiliary tasks, such as manipulation type identification [39, 40], manipulation parameter estimation [41], blending boundary detection [23], and face reconstruction [24, 27]), leading to a multitask learning paradigm [22, 23, 42, 24, 25, 26, 27, 28, 43]. Here, the underlying assumption is that the primary task will likely gain from auxiliary tasks through knowledge transfer. However, most auxiliary tasks are specific to manipulations, which may hinder the generalizability of DeepFake detectors. Recently, Zou et al. [44] introduced a semantics-oriented multitask learning paradigm for DeepFake detection. They argued that face forgery stems from computational methods that modify semantic face attributes beyond the thresholds of human discrimination. Although conceptually and computationally appealing, training semantics-oriented DeepFake detectors would necessitate considerable human annotations to specify manipulation degree parameters and semantic label relationships [44]. Moreover, the allocation of task-agnostic and task-specific parameters is optimized manually, which is bound to be suboptimal in facilitating knowledge transfer among tasks.
Figure 1: Illustration of the label hierarchy in our expanded FF++ dataset [20]. The hierarchy organizes manipulation types into three nodes: identity (face-swapping via Deepfakes [1] and FaceSwap [45]), expression (mouth editing via Face2Face [46] and NeuralTextures [4]), and physical_inconsistency (local face editing via data augmentations). These nodes connect to six leaf nodes representing six distinct local face regions. Edges denote manipulation relationships: identity affects all regions, expression targets lip and mouth, while physical_inconsistency modifies eye, lip, mouth, and nose.
@@
In this paper, we further pursue the promising direction of semantics-oriented multitask learning for DeepFake detection with two important improvements. First, we present a dataset expansion technique that automatically enriches the binary labeling of existing DeepFake datasets to semantic hierarchical labeling while increasing their sizes through data augmentation (see Fig. 1). A crucial element is to incorporate the physical_inconsistency node that encompasses local face region manipulations, without manually identifying interdependent changes in other global face attributes (if any). Second, unlike discriminative learning that directly predicts labels from images, we take inspiration from the recent advances in vision-language correspondence [47] and encode both the input image and label hierarchy (represented by text templates) in a common feature space. This enables us to view model parameter sharing/splitting in multitask learning from the perspective of model capacity allocation. More specifically, we share the image encoder parameters across all tasks, while the capacity allocated to each individual task is end-to-end optimized. In addition, we adopt bi-level optimization [48] to prioritize the primary task by minimizing its fidelity loss [49] with respect to the loss weighting vector. The parameters of the DeepFake detector are trained by minimizing the weighted sum of fidelity losses of individual tasks.
@@
Extensive experiments on six datasets [20, 50, 51, 52, 53, 44] demonstrate that our Semantics-oriented Joint Embedding DeepFake Detector (SJEDD) consistently improves the cross-dataset, cross-manipulation, and cross-attribute detection performance, compared to state-of-the-art DeepFake detectors. In summary, our contributions are threefold:
  • an automated dataset expansion technique that enriches the binary labeling of DeepFake datasets,
  • a semantics-oriented joint embedding method for DeepFake detection, coupled with bi-level optimization to fully automate the training process, and
  • an experimental demonstration that the proposed SJEDD offers better generalizability and interpretability.
@@ II Related Work
In this section, we provide a concise review of DeepFake detectors, emphasizing the multitask learning paradigm. We also discuss joint embedding architectures in the broader context of machine learning.
@@ II-A DeepFake Detectors
Non-learning-based detectors rely on extensive domain expertise. The first type of knowledge involves mathematical models of the image source, which summarizes the essential elements of what constitutes an authentic image [11]. The second type of knowledge involves computational models of the digital photography pipeline, along which each step can be exploited to expose forgery [54], including the lens distortion and flare [12, 13], color filter array [55], sensor noise [14, 56], pixel response non-uniformity [57], camera response function [58], and file storage format [59]. The third type of knowledge involves physical and geometric models of our three-dimensional world, especially those that capture the interactions of light, optics, and objects. Face forgery can be exposed by detecting physical and geometric inconsistencies, such as those found in illumination [15], shadow [16], reflection [17], and motion [60]. Another line of research that has received significant attention involves detecting abnormal physiological reactions from eye blinking [61], pupil shape [62], head pose [63], and corneal specularity [64]. Nevertheless, these physiological features may not sufficiently capture the nuances needed to identify face forgery.
@@
Learning-based methods learn to detect DeepFake images and videos from data. Afchar et al. [19] proposed to examine DeepFake videos at the mesoscopic level. Zhou et al. [18] designed a two-stream network to jointly analyze camera characteristics and visual artifacts. Li and Lyu [65] aimed to detect affine warping artifacts, while Li et al. [23] and Nguyen et al. [66] focused on locating image blending boundaries. Dong et al. [26] revealed a shortcut termed “implicit identity leakage” and proposed an identity-unaware DeepFake detector. Gu et al. [67] and Zheng et al. [68] examined spatiotemporal inconsistencies at the signal level, while Haliassos et al. [29] did so at the semantic level through lipreading pretraining. Xu et al. [69] proposed a thumbnail layout to trade spatial modeling for temporal modeling. Additionally, complex architectures such as capsule networks [70] and vision Transformers [71, 32, 72], along with advanced training strategies such as adversarial training [25, 73], test-time training [74], graph learning [34, 36], contrastive learning [33, 38], and one-class classification [75], have also been explored for DeepFake detection. Generally, learning-based methods typically overfit to training-specific artifacts [18, 65, 26, 56, 23, 76].
@@
Multitask learning-based methods have emerged as a promising offshoot of learning-based methods for DeepFake detection in the past five years. By leveraging auxiliary supervisory signals during training, these methods tend to offer improved generalizability compared to single-task approaches. Notable auxiliary tasks include manipulation localization [22, 25] (with blending boundary detection [23] as a special case), manipulation classification and attribution [39, 40], manipulation parameter estimation [41], face image reconstruction [24, 27], and tasks that promote robust and discriminative feature learning [28]. Most auxiliary tasks for DeepFake detection are manipulation-oriented, which may overfit detectors to manipulation-specific patterns and limit their effectiveness against sophisticated, novel forgery. To address this limitation, Zou et al. [44] pioneered a paradigm shift from manipulation-oriented to semantics-oriented detection. Their approach emphasizes the joint learning of high-level representations of global face attributes and low-level representations of local face regions, thus enhancing detection generalizability. In this paper, we further advance this promising paradigm. From the data perspective, we introduce an automated dataset expansion technique that not only enlarges existing DeepFake datasets but also enriches them with a collection of labels organized in a semantic hierarchical graph (see Fig. 1). From the model perspective, we leverage joint embedding instead of discriminative learning to automate model parameter sharing/splitting and to facilitate task relationship modeling. Concurrently, Sun et al. [35] proposed VLFFD, which also implemented DeepFake detection via joint embedding and multitask learning. At a high level, SJEDD employs semantics-oriented multitask learning, while VLFFD adopts a manipulation-oriented approach, centered on testing the feasibility of CLIP [47] for DeepFake detection. This distinction drives four key computational differences. First, SJEDD uses probabilistic modeling on a semantic hierarchical graph, while VLFFD relies on the standard CLIP architecture. Second, SJEDD prioritizes semantics-oriented tasks with text prompts, whereas VLFFD focuses on identifying manipulation artifacts. Third, SJEDD applies principled bi-level optimization for automated multitask learning, contrasting with the heuristic tuning of task-specific hyperparameters in VLFFD.
@@
Figure 2: Pipeline for local face region manipulation. When searching for the best candidate as the source image, we minimize the Euclidean distance between the $68$ detected landmarks [23] of the target and candidate face images (excluding those with the same identity).
@@ II-B Joint Embedding Architectures
Joint embedding architectures [77] aim to learn embeddings that are close for compatible inputs (e.g., an input image and its target label) and distant for incompatible inputs, as opposed to discriminative approaches that directly predict labels from inputs. This idea can be traced back to Siamese networks [78], initially proposed for signature verification and later adapted to other computer vision tasks like face verification [79] and recognition [80], and video feature learning [81]. With the rise of self-supervised learning, a thorough exploration of joint embedding architectures is conducted, contrastively (e.g., MoCo [82] and SimCLR [83]) and non-contrastively (e.g., Barlow Twins [84] and I-JEPA [85]). The more recent vision-language models [86, 47] are also manifestations of joint embedding architectures and are quickly adapted to multiple downstream tasks such as image quality assessment [87], realistic image enhancement [88], image-text retrieval [89], and multimodal fake news detection [90].
Recent DeepFake detectors have also leveraged vision-language joint embedding [91, 92, 35, 93, 94, 95]. While these methods directly utilize cosine similarity scores for detection, the proposed SJEDD maps these scores into a joint probability distribution over a customized semantic label hierarchy. Our design explicitly integrates label relationships and unifies low-level visual features with high-level semantic cues, thus facilitating semantics-oriented DeepFake detection.
@@ III Semantics-Oriented Dataset Expansion: A Demonstration on FF++
In this section, we describe our automated dataset expansion technique, consisting of two steps: data augmentation and data relabeling. Data augmentation generates new images by manipulating local face regions. Data relabeling automatically assigns DeepFake images with a set of labels through semantic contextualization of the adopted face manipulations [44]. To demonstrate the proposed dataset expansion technique, we apply it to expand the widely used FF++ dataset [20].
@@ III-A Data Augmentation
We augment FF++ by manipulating four local face regions, i.e., eye, lip, mouth, and nose, as illustrated in Fig. 2. Given a real face as the target, we initially search for the best source face (real or fake) by minimizing the Euclidean distance between their corresponding $68$ detected landmarks [96] while excluding faces with the same identity [23]. Next, we linearly blend the local face region from the source, delineated by a Gaussian-blurred mask, into the target image, followed by color correction [97]. In our implementation, candidate face images are sourced from the Deepfakes [1] and FaceSwap [45] subsets of FF++. For each of the four local face regions, the number of augmented images matches that of forged images by each face manipulation method in the original FF++.
TABLE I: Semantical hierarchical labeling of the expanded FF++ dataset: “1” indicates that the face attribute (or region) has been modified, “0” means no modification, and “N/A” denotes that the underlying label is unobserved
Manipulation / Region Face Expression Identity
Physical
Inconsistency
Eye Eyebrow Lip Mouth Nose Skin
@@
III-B Data Relabeling
Inspired by [44], we first construct a semantic hierarchical graph to contextualize the four face manipulations of FF++: Deepfakes [1], FaceSwap [45], Face2Face [46], and NeuralTextures [4]. The former two swap faces between images to achieve identity manipulation, while the latter two change the expression of target faces through manipulation of the mouth region. As a result, we instantiate two global face attribute nodes: expression and identity. We also include six leaf nodes, representing nonoverlapping local face regions: eye, eyebrow, lip, mouth, nose, and skin. For the two identity manipulations based on face swapping, all local face regions are assumed to be altered. In contrast, for the two expression manipulations, only lip and mouth regions are affected. To accommodate local face manipulations in the graph, we introduce an additional global face attribute node, physical_inconsistency, which signifies images where local face regions are replaced by the corresponding regions from source images, resulting in physical incompatibility.
@@ III-C Discussion
The proposed dataset expansion technique allows some labels in the semantic hierarchical graph to remain unobserved, which improves labeling reliability. For instance, in the case of face swapping, it is unclear whether the expression has been altered, so we leave this label unobserved. Meanwhile, in the expanded dataset, a single face manipulation can simultaneously alter multiple face attributes (e.g., identity and physical_consistency by FaceSwap [45]), while different manipulations can also modify the same attribute (e.g., identity by Deepfakes [1] and FaceSwap). As such, our technique emphasizes a semantics-oriented rather than manipulation-oriented contextualization of face manipulations akin to FFSC [44] and achieves so without expensive human labeling. Table I presents the list of face manipulations in the expanded FF++ dataset, along with the corresponding modified face attributes and regions.